# Quantum information with quantum-like bits

### v1

@@ I Introduction
In the spirit of quantum-inspired computing, we showed in recent work that a large class of synchronizing classical systems can be utilized as a resource for information processing. Scholes (2020, 2023, 2024); Scholes and Amati (2024) The robustness of these systems to noise and stochastic disorder improves with size, following a somehow opposite and favorable trend compared to genuine quantum resources. This well-known behavior is related to the occurrence of classical emergent phenomena and synchronization. Haken (1975); Fano (1992); Strogatz (2012); Artime and De Domenico (2022) Graph theory is the natural mathematical framework used to describe synchronizing many-body systems. Scholes (2020, 2023); Kassabov, Strogatz, and Townsend (2022); Townsend, Stillman, and Strogatz (2020); Rodrigues et al. (2016) In particular, emergent states can be clearly distinguished from a broad band of random states in the spectrum of the adjacency matrices of Erdös-Renyi random graphs and expander families, such as the $k$-regular graphs. Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) In Ref. Scholes, 2024, we mapped emergent states onto a computational basis of quantum-like (QL) bits, which can be easily identified from sharp and isolated peaks in the spectra of regular graphs. This feature implies that QL states are easily distinguishable and measurable even in presence of noise and dissipation, that can usually broaden the spectral lines of the network. In Ref. Scholes and Amati, 2024 we built on these results and proposed a general structure for the resources of multiple QL bits, based on the Cartesian products of one-QL-bit graphs.
Our approach is derived consistently with a QL probabilistic framework proposed by Khrennikov. Khrennikov (2007a, 2019, 2010) This theory shows that genuine quantum systems are not strictly required to process information in a QL way. Instead, to observe QL phenomena it is sufficient to construct a contextual probability theory involving suitable notions of QL states, observables, and their measurement.
In the present paper we extend the foundational work of Ref. Scholes, 2024; Scholes and Amati, 2024; Scholes, 2023 to formalize the notion of QL states and gates. Furthermore, consistently with Khrennikov’s approach, Khrennikov (2010) we list a set of measurement axioms which allows to define a generalized notion of nonclassical interference between observable operators.
@@
The present paper is structured as follows. In ?? we summarize the formalism describing how to construct a single-QL-bit Hilbert space. In ?? we discuss its generalization to higher dimensions, and we show explicitly in ?? how to construct Bell states for two QL bits. In ?? we study the action of a complete set of QL gates, and we provide concrete examples of their action. Finally, in ??, we present a set of axioms that enables us to define a QL probabilistic theory in our approach. In particular, we highlight the presence of a nonclassical interference term in the relation between total and conditional probabilities, specific to non-Kolmogorovian contextual probabilistic models. Khrennikov (2010)
@@ II Quantum-like bits
In this section we derive a mapping a from classical network to the Hilbert space of interacting QL bits. While the concept has been already elucidated in Ref. Scholes, 2024, here we lay out a more formal description. Later in the paper we use this formalism to study the notion of gates to perform quantum information processing, and we provide concrete examples.
@@ II.1 Isolated system
Let us consider two classical many-body systems with pairwise interactions between the elementary constituents. Those systems can be mapped onto two graphs $G_{1}$ and $G_{2}$. We assume here that both graphs involve the same number of vertices ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices. For simplicity we focus here on undirected simple graphs, without loops. The adjacency matrix for this class of graphs is symmetric with entries either 00 or $1$ and null diagonal. Additionally, we assume an isotropic structure for all two-body interactions, such that the valence of the graphs is constant and equal to $k$. Diestel (2017) In particular,
@@
involving a coupling matrix $C$. Note that $C$ does not represent a simple graph by itself, i.e. it is in general non-symmetric, and it can admit diagonal entries mapping corresponding vertices in $A_{1}$ and $A_{2}$. Conversely, ??, being overall symmetric and loop-free, describes a simple undirected graph. We sample $C$ by randomly assigning $1$ to $N_{\mathrm{C}}\leq{N_{G}}^{2}$ of its entries. This matrix is chosen to be regular with valency $l=N_{\mathrm{C}}/{N_{G}}$ [similarly to ??]. ?? represents the fundamental resource used to build an isolated QL bit. Scholes (2024)
@@
As discussed in Refs. Scholes, 2023, 2024, resources with structure as in ?? are particularly appealing for information processing. In fact, these resources allow for emergent states associated to distinguishable eigenvalues, well isolated from a large band of random states. By mapping the states associated to those peaks onto a Hilbert space, we obtain a platform for information logic that can be controlled by network manipulation. We will elaborate on this important concept further in the manuscript.
@@
$\psi_{+}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}\mapsto\ket{+},\hskip 17.07164pt\psi_{-}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\\ -a_{2}\end{pmatrix}\mapsto\ket{-},$ (II.4)
where $a_{i}$ is the normalized eigenvector associated to the largest eigenvalue of $A_{i}$. The vectors $\psi_{\sigma}$, $\sigma\in\{+,-\}$, are eigenvectors of ?? with eigenvalues $k\pm l$. In the following we show that the mapping in ?? leads to a well-defined Hilbert space. First of all, it is straightforward to show that the $\psi_{\sigma}$’s are mutually orthogonal and normalized:
@@
where $\sigma,\sigma^{\prime}\in\{+,-\}\equiv\mathcal{S}$. Operators acting on ?? are expressed in terms of the matrices in $M_{2{N_{G}}}(\mathbb{R})$
@@ II.2.1 Formalism
where each $\mathcal{R}^{(q)}$ has the structure shown in ??. The tensors
$\psi_{\bm{\sigma}}=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(q)}}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{\sigma^{(q)}}=\ket{\bm{\sigma}},$ (II.9)
@@
$[\psi_{\bm{\sigma}}]_{\bm{i}}$ $=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}.$ (II.10)
As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\{+,-\}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$. Let us label this eigenvalue $\lambda_{\sigma^{(q)}}$. The action of $\mathcal{R}$ on $\psi_{\bm{\sigma}}$ is then given by
@@
$=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}\psi_{\bm{\sigma}}=\lambda_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (II.11)
@@
$\lambda_{\bm{\sigma}}=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}.$ (II.12)
@@
$=\prod_{q=1}^{{N_{\mathrm{QL}}}}\delta_{{\sigma}^{(q)},{\sigma^{\prime}}^{(q)}}=\prod_{q=1}^{{N_{\mathrm{QL}}}}\braket{\sigma^{(q)}}{{\sigma^{\prime}}^{(q)}}.$ (II.13)
@@
$\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}$ $=\psi_{\bm{\sigma}}\otimes\psi_{\bm{\sigma}^{\prime}}^{\dagger}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{{\sigma}_{q}^{(q)}}\bra{{\sigma^{\prime}}_{q}^{(q)}}=\ket{\bm{\sigma}}\bra{\bm{\sigma}^{\prime}},$ (II.14)
@@
$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}\right]_{\bm{i,j}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}\psi^{*}_{\sigma^{\prime(q)},j^{(q)}}.$ (II.15)
@@
$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}^{\dagger}\right]_{\bm{ij}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},j^{(q)}}^{*}\psi_{\sigma^{\prime(q)},i^{(q)}}=\left[\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right]_{\bm{ij}},$ (II.16)
@@
$\Psi_{\bm{\sigma\sigma^{\prime}}}\Psi_{\bm{\pi\pi^{\prime}}}=\Psi_{\bm{\sigma\pi^{\prime}}}\braket{\bm{\psi_{\sigma^{\prime}}},\bm{\psi_{\pi}}}.$ (II.17)
@@ II.2.2 Example: Bell states
$\mathcal{R}=\mathcal{R}^{(1)}\otimes\mathbb{1}_{2{N_{G}}}+\mathbb{1}_{2{N_{G}}}\otimes\mathcal{R}^{(2)}.$ (II.18)
@@
$\rho(\lambda)=\int_{-\infty}^{+\infty}\mathrm{d}x\;\rho^{(1)}(x)\rho^{(2)}(\lambda-x),$ (II.19)
where $\rho^{(q)}(x)$ is the spectrum of a single QL bit, analogously to what shown in ??. The generalization of ?? provides a simple recipe to construct the spectrum of a system of an arbitrary number of QL bits starting from the spectra for ${N_{\mathrm{QL}}}=1$. The distribution calculated from the convolution (conv.) ??, is shown as a green (light) dashed curve in ??, and it agrees satisfactorily with the direct calculation. We mark in red on the picture the isolated peaks corresponding to the emergent eigenvalues of the two-QL-bit systems. These are $\lambda_{++}=2k+2l$ (circle), $\lambda_{+-}=\lambda_{-+}=2k$ (cross) and $\lambda_{--}=2k-2l$ (square). These eigenvalues are associated to the corresponding eigenvectors $\{\psi_{\sigma\sigma^{\prime}}\}_{\sigma,\sigma^{\prime}\in\mathcal{S}}$, defining an orthonormal basis for the Hilbert space of two QL bits.
@@
$\psi^{(1)}_{\pm}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{\pm}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.20)
where $a_{i}$’s and $b_{i}$’s are defined as in ?? for the resources ${\mathcal{R}}^{(1)}$ and ${\mathcal{R}}^{(2)}$, respectively. With ?? we can express
$\psi_{+\pm}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\pm a_{2}\otimes b_{2}\end{pmatrix},$ (II.21a)
$\psi_{-\pm}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}&\mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.21b)
@@
$\Phi^{\mathrm{Bell}}_{\pm}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{++}\pm\psi_{--}\right),$ (II.22a)
$\Psi^{\mathrm{Bell}}_{\pm}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{+-}\pm\psi_{-+}\right),$ (II.22b)
@@
$\Phi^{\mathrm{Bell}}_{+}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&a_{2}\otimes b_{2}\end{pmatrix},$ (II.23a)
$\Phi^{\mathrm{Bell}}_{-}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix},$ (II.23b)
$\Psi^{\mathrm{Bell}}_{+}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&-a_{2}\otimes b_{2}\end{pmatrix},$ (II.23c)
$\Psi^{\mathrm{Bell}}_{-}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&-a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix}.$ (II.23d)

### v4

@@ I Introduction
In the spirit of quantum-inspired computing, we showed in recent work that a large class of synchronizing classical systems can be utilized as a resource for information processing. Scholes (2020, 2023, 2024); Scholes and Amati (2024) We focused, in particular, on many body systems whose correlations obey the graph structure shown in panel a) of ??. Here, two densely connected regions (in blue) of a network are mutually interacting via sparse links (in red). A special feature of this class of networks is that they produce pairs of emergent states, obtained by the direct sum of the emergent states of the blue subgraphs, with identical and opposite phases. The key requirement of the blue subgraphs, then, is that they exhibit an emergent state—a common feature of Erdös-Renyi random graphs and expander families, such as the k-regular graphs. Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) In our framework, emergent states can be clearly distinguished from a broad band of states arising from random network configurations (dashed lines in the spectra of ??). Their robustness against noise and stochastic disorder increases with system size, exhibiting a favorable trend in contrast to the typical behavior of genuine quantum resources. Scholes (2023, 2020) This intriguing feature is connected to the occurrence of classical synchronization phenomena. Haken (1975); Fano (1992); Strogatz (2012); Artime and De Domenico (2022) We showed in Ref. Scholes, 2024 that the emergent states of connected expander subgraphs are isomorphic to the Hilbert space of a qubit. This led us to introducing the notion of quantum-like (QL) bits within the context of classically-synchronizing networks. In Ref. Scholes and Amati, 2024 we built on these results and proposed a general framework to construct resources for multiple QL bits. In particular, we showed that the Cartesian product ($\square$) of ${N_{\mathrm{QL}}}$ bipartite graphs obeys the same dimensionality of a register of ${N_{\mathrm{QL}}}$ qubits [see panel b) of ??].
In the present paper, we build on the foundational work of Ref. Scholes, 2024; Scholes and Amati, 2024; Scholes, 2023 to formalize the concept of QL states and gates. We present a computational framework for implementing information processing within the QL formalism through appropriate manipulations of the emergent states of network resources. As depicted in panel c) of ??, we show that an arbitrary gate transformation $V_{\mathrm{g}}$ can be expressed both as graph transformation, and as a map of the emergent states of the graph. Finally, we propose a measurement approach upon information processing that aligns with the QL probabilistic framework developed by Khrennikov Khrennikov (2007a, 2019, 2010). This theory shows that genuine quantum systems are not strictly required to process information in a QL way. Instead, to observe QL phenomena it is sufficient to construct a contextual probability theory involving suitable notions of QL states, observables, and their measurement. Consistently with Khrennikov’s approach, Khrennikov (2010) we list a set of measurement axioms which allows to define a generalized notion of nonclassical interference between observable operators.
@@
The present paper is structured as follows. In ?? we summarize the formalism describing how to construct a single-QL-bit Hilbert space. In ?? we discuss its generalization to higher dimensions, and we show explicitly in ?? how to construct Bell states for two QL bits. In ?? we study the action of a complete set of QL gates, and we provide concrete examples of their action. In ??, we present a set of axioms that enables us to define a QL probabilistic theory in our approach. In particular, we highlight the presence of a nonclassical interference term in the relation between total and conditional probabilities, specific to non-Kolmogorovian contextual probabilistic models. Khrennikov (2010) Finally, in ??, we outline the complete workflow for QL information processing and discuss its efficiency and accuracy.
@@
Figure 1: Illustration of the main concepts presented in this paper and in prior work by the authors. Scholes (2024); Scholes and Amati (2024) a) Many body systems consisting of networks of synchronizing dipoles can easily exhibit classical synchronization phenomena. By structuring two-body correlations into two densely connected regions (blue) weakly interacting through sparse links (red), synchronization gives rise to pairs of emergent states. These states are well-distinguishable from a large band of random states at lower energies (dashed line in the sketch of the spectrum). In Ref. Scholes, 2024 we proposed a map from emergent states to the computational basis of a qubit. b) In Ref. Scholes and Amati, 2024, we demonstrated that the emergent states of the Cartesian product of a set of synchronizing graphs scale with the same dimensionality of a qubit system. In this paper, we expand over these concepts by providing further mathematical details on the formalism for this many-body problem. c) Additionally, we elaborate on the concept of information processing within the QL framework. We show that operations on graphs, implemented through suitable transformations of their edges, are equivalent to gate transformations $V_{\mathrm{g}}$ applied directly to the emergent states.
@@ II Quantum-like bits
In this section we derive a mapping from a classical network to the Hilbert space of interacting QL bits. While the concept has been already elucidated in Ref. Scholes, 2024, here we lay out a more formal description. Later in the paper we use this formalism to study the notion of gates to perform quantum information processing, and we provide concrete examples.
@@ II.1 Isolated system
Let us consider two classical many-body systems with pairwise interactions between the elementary constituents. Those systems can be mapped onto two graphs $G_{1}$ and $G_{2}$. We assume here that both graphs involve the same number of vertices, ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices. For simplicity we focus here on undirected simple graphs, without loops. The adjacency matrix for this class of graphs is symmetric with entries either 00 or $1$ and null diagonal. Additionally, we assume an isotropic structure for all two-body interactions, such that the valence of the graphs is constant and equal to $k$. Diestel (2017) In particular,
@@
involving a coupling matrix $C$. Note that $C$ does not represent a simple graph by itself, i.e. it is in general non-symmetric, and it can admit diagonal entries connecting corresponding vertices in $A_{1}$ and $A_{2}$. Conversely, ??, being overall symmetric and loop-free, describes a simple undirected graph. We sample $C$ by randomly assigning $1$ to $N_{\mathrm{C}}\leq{N_{G}}^{2}$ of its entries. This matrix is chosen to be regular with valency $l=N_{\mathrm{C}}/{N_{G}}$ [similarly to ??]. ?? represents the fundamental resource used to build an isolated QL bit. Scholes (2024)
@@
As discussed in Refs. Scholes, 2023, 2024, resources with structure as in ?? are particularly appealing for information processing. In fact, these resources allow for emergent states associated to distinguishable eigenvalues, well isolated from a large band of random states. By mapping the states associated to those peaks onto a Hilbert space, we obtain a platform for information processing that can be controlled by network manipulation. We will elaborate on this important concept further in the manuscript.
@@
$\psi_{\downarrow}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\\ -a_{2}\end{pmatrix}\mapsto\ket{\downarrow},\hskip 17.07164pt\psi_{\uparrow}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}\mapsto\ket{\uparrow}$ (II.4)
where $a_{i}$ is the normalized eigenvector associated to the largest eigenvalue of $A_{i}$. The vectors $\psi_{\sigma}$, $\sigma\in\{\downarrow,\uparrow\}\equiv\mathcal{S}$, are eigenvectors of ?? with eigenvalues $\{k-l,k+l\}$, respectively. The current arrow notation for the computational basis helps prevent any ambiguity with additional numeric subscripts and superscripts, which are used here and in the following sections for different purposes. In prior work we used the intention of the QL bit construction to represent explicitly the basis of the the subgraphs, so that the emergent state $a_{1}$ of the graph with adjacency matrix $A_{1}$ represents $\ket{\downarrow}$, and similarly $a_{2}$ represents $\ket{\uparrow}$.Scholes (2024); Scholes and Amati (2024) ?? describes another basis choice, equally valid. The present approach establishes a direct connection between the computational basis and the emergent states of the composite graph described by ??. In the following we show that the mapping in ?? leads to a well-defined Hilbert space. First of all, it is straightforward to show that the $\psi_{\sigma}$’s are mutually orthogonal and normalized:
@@
where $\sigma,\sigma^{\prime}\in\mathcal{S}$. Operators acting on ?? are expressed in terms of the matrices in $M_{2{N_{G}}}(\mathbb{R})$
@@ II.2.1 Formalism
where each $\mathcal{R}^{(q)}$ has the structure shown in ??. ?? is the adjacency matrix of the Cartesian product graph Scholes and Amati (2024)
[…]
Here, $\mathcal{G}^{(q)}$ is the single-QL-bit graph with adjacency matrix $\mathcal{R}^{(q)}$. The tensors
$\psi_{\bm{\sigma}}=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(q)}}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{\sigma^{(q)}}=\ket{\bm{\sigma}},$ (II.10)
@@
$[\psi_{\bm{\sigma}}]_{\bm{i}}$ $=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}.$ (II.11)
As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\{0,1\}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$. Let us label this eigenvalue $\lambda_{\sigma^{(q)}}$. The action of $\mathcal{R}$ on $\psi_{\bm{\sigma}}$ is then given by
@@
$=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}\psi_{\bm{\sigma}}=\lambda_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (II.12)
@@
$\lambda_{\bm{\sigma}}=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}.$ (II.13)
@@
$=\prod_{q=1}^{{N_{\mathrm{QL}}}}\delta_{{\sigma}^{(q)},{\sigma^{\prime}}^{(q)}}=\prod_{q=1}^{{N_{\mathrm{QL}}}}\braket{\sigma^{(q)}}{{\sigma^{\prime}}^{(q)}}.$ (II.14)
@@
$\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}$ $=\psi_{\bm{\sigma}}\otimes\psi_{\bm{\sigma}^{\prime}}^{\dagger}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{{\sigma}_{q}^{(q)}}\bra{{\sigma^{\prime}}_{q}^{(q)}}=\ket{\bm{\sigma}}\bra{\bm{\sigma}^{\prime}},$ (II.15)
@@
$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}\right]_{\bm{i,j}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}\psi^{*}_{\sigma^{\prime(q)},j^{(q)}}.$ (II.16)
@@
$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}^{\dagger}\right]_{\bm{ij}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},j^{(q)}}^{*}\psi_{\sigma^{\prime(q)},i^{(q)}}=\left[\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right]_{\bm{ij}},$ (II.17)
@@
$\Psi_{\bm{\sigma\sigma^{\prime}}}\Psi_{\bm{\pi\pi^{\prime}}}=\Psi_{\bm{\sigma\pi^{\prime}}}\braket{\bm{\psi_{\sigma^{\prime}}},\bm{\psi_{\pi}}}.$ (II.18)
@@ II.2.2 Example: Bell states
$\mathcal{R}=\mathcal{R}^{(1)}\otimes\mathbb{1}_{2{N_{G}}}+\mathbb{1}_{2{N_{G}}}\otimes\mathcal{R}^{(2)}.$ (II.19)
@@
$\rho(\lambda)=\int_{-\infty}^{+\infty}\mathrm{d}x\;\rho^{(1)}(x)\rho^{(2)}(\lambda-x),$ (II.20)
where $\rho^{(q)}(x)$ is the spectrum of a single QL bit, shown in ??. The generalization of ?? provides a simple recipe to construct the spectrum of a system of an arbitrary number of QL bits starting from the spectra for ${N_{\mathrm{QL}}}=1$. The distribution calculated from the convolution (conv.) ??, is shown as a green (light) dashed curve in ??, and it agrees satisfactorily with the direct calculation. We mark in red on the picture the isolated peaks corresponding to the emergent eigenvalues of the two-QL-bit systems. These are $\lambda_{11}=2k+2l$ (circle), $\lambda_{10}=\lambda_{01}=2k$ (cross) and $\lambda_{00}=2k-2l$ (square). These eigenvalues are associated to the corresponding eigenvectors $\{\psi_{\sigma\sigma^{\prime}}\}_{\sigma,\sigma^{\prime}\in\mathcal{S}}$, defining an orthonormal basis for the Hilbert space of two QL bits.
@@
$\psi^{(1)}_{[\uparrow\downarrow]}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{[\uparrow\downarrow]}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.21)
where $a_{i}$’s and $b_{i}$’s are defined as in ?? for the resources ${\mathcal{R}}^{(1)}$ and ${\mathcal{R}}^{(2)}$, respectively. The square brackets indicate that the sign on the right-hand side of each expression (+ or -) is chosen depending on the direction of the arrow on the left-hand side ($\uparrow$ or $\downarrow$, respectively). With ?? we can express
$\psi_{\uparrow,[\uparrow\downarrow]}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\pm a_{2}\otimes b_{2}\end{pmatrix},$ (II.22a)
$\psi_{\downarrow,[\uparrow\downarrow]}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}&\mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.22b)
@@
$\Phi^{\mathrm{Bell}}_{[\uparrow\downarrow]}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\uparrow}\pm\psi_{\downarrow\downarrow}\right),$ (II.23a)
$\Psi^{\mathrm{Bell}}_{[\uparrow\downarrow]}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\downarrow}\pm\psi_{\downarrow\uparrow}\right),$ (II.23b)
@@
$\Phi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&a_{2}\otimes b_{2}\end{pmatrix},$ (II.24a)
$\Phi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix},$ (II.24b)
$\Psi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&-a_{2}\otimes b_{2}\end{pmatrix},$ (II.24c)
$\Psi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&-a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix}.$ (II.24d)